Extended Poincaré parasupergroup with central charges

Nikitin A.G.

Institute of Mathematics of tha Academy of Sciences of Ukraine, Kiev, Ukraine

Irreducible representations of the extended Poincaré parasuperalgebra with non-trivial central charges are described. Linear and non-linear models invariant w.r.t. the related parasupergroup are discussed.